If we require insertionSort() to sort a list of 10-billion elements in no more than one minute, how many times faster would your new computer system need to be than the fastest system currently in existence, i.e., the one that performs a comparison in 25 ns?
Answer #1

In insertion sort, we know that in worst case,there can be n(n-1)/2 comparisons and n(n-1)/2 swaps.

From above given information I am assuming that one comparison one operation takes 25ns in fastest computer available.

Assuming fastest computer takes 25ns in both operations(i.e comparison and swaps)

Total time taken by insertion sort in fastest computer= Total number of comparisons * time taken by single comparison + Total number of swaps*time taken by single swaps

=( n(n-1)/2)*25ns + (n(n-1)/2)*25ns

= 25n(n-1)

Now put n = 10 billion = 10 * 109 = 1010

Time taken by insertion sort in fastest computer = 1010(1010-1)*25 * 10-9

= (1020-1010)*25*10-9

= 25 *(1011-10)

= (25 * 1011 - 250)seconds

Ignore -250 in above expression because it would not affect our answer as 25*1011 is very very big number compared to 250.

So,time taken by insertion sort in fastest computer in existence= 25*1011s

Now,we have a computer system in which we want time taken by insertion sort <=60 seconds

Do, required speedup = execution time in fastest computer / execution time required in new computer system

= 25*1011/ 60

= 0.4167 * 1011

= 4.167 * 1010

Required speedup = 4.167 * 1010

​​​so,in order for new system to execute insertion sort on 10 billion elements in no more than 1 minute it needs to 4.167 * 1010 faster than current fastest available system.